# Autoencoding Random Forests

### v1

@@ Appendix C Leaf assignments via lasso and greedy search
While the algorithm is guaranteed to converge on a consistent set of leaf assignments, this may only be achievable by random sampling of leaves under a consistency constraint. Such uninformative results are likely the product of noisy estimates for $\hat{\mathbf{K}}_{0}$. Note that while the maximal clique problem is famously NP-complete [56], we must solve this combinatorial task at most once (at $t=1$). In subsequent rounds $t\geq 2$, we simply check whether any new nodes have been added to $C(t-1)$. Maximal clique solvers are highly optimized and have worst-case complexity $\mathcal{O}(2^{B/3})$, but are often far more efficient due to clever heuristics [96]. In practice, the method tends to find reasonable leaf assignments in just a few rounds. When computing leaf assignments for multiple test points, we may run Alg. 1 in parallel, as solutions are independent for each sample.
@@
[…]
@@
$q^{(b)}(t)\leftarrow\operatorname*{arg\,max}_{i\in[d_{\Phi}^{(b)}]}\leavevmode\nobreak\ \hat{p}_{i}^{(b)}\quad\leavevmode\nobreak\ \text{s.t.}\leavevmode\nobreak\ \leavevmode\nobreak\ R_{i}^{(b)}\cap S(t)\neq\emptyset$
@@ Appendix D Computational complexity
The encoding step requires $\mathcal{O}(n^{2})$ space to store the adjacency matrix. For large datasets, this can be approximated by using a subsample of all training points to compute $\mathbf{K}$, at the cost of losing the exact functional equivalence of Eq. 2. Similar tricks are common in kernel methods such as Gaussian process regression [72].
@@
The split relabeling method is essentially an iterated CART algorithm, and therefore requires $\mathcal{O}(d_{\mathcal{Z}}\tilde{n}\log\tilde{n})$ time to relabel each split, where $\tilde{n}$ is the number of synthetic samples $\tilde{\mathbf{X}}$. However, since the number of splits is generally exponential in tree depth, this can still be challenging for deep forests.
The most efficient decoder is the $k$-NN method, which proceeds in two steps. First, we find the nearest neighbors in embedding space, which for reasonably small $d_{\mathcal{Z}}$ can be completed in $\mathcal{O}(m\log n)$ time using kd-trees. Next, we compute the intersection of all leaf regions for selected neighbors, incurring a cost of $\mathcal{O}(kmd_{\Phi}d_{X})$.

### v2

@@ Appendix C Leaf assignments via lasso and greedy search
While the algorithm is guaranteed to converge on a consistent set of leaf assignments, this may only be achievable by random sampling of leaves under a consistency constraint. Such uninformative results are likely the product of noisy estimates for $\hat{\mathbf{K}}_{0}$. Note that while the maximal clique problem is famously NP-complete [61], we must solve this combinatorial task at most once (at $t=1$). In subsequent rounds $t\geq 2$, we simply check whether any new nodes have been added to C(t1)C(t-1). Maximal clique solvers are highly optimized and have worst-case complexity $\mathcal{O}(2^{B/3})$, but are often far more efficient due to clever heuristics [102]. In practice, the method tends to find reasonable leaf assignments in just a few rounds. When computing leaf assignments for multiple test points, we may run Alg. 1 in parallel, as solutions are independent for each sample.
@@
1:  Initialize: $t\leftarrow 0$, $C(t)\leftarrow\emptyset$, $S(t)\leftarrow\mathcal{X}$, $\texttt{converged}\leftarrow\texttt{FALSE}$
@@
$q^{(b)}(t)\leftarrow\operatorname*{arg\,max}_{i\in[d_{\Phi}^{(b)}]}\penalty 10000\ \hat{p}_{i}^{(b)}\qquad\penalty 10000\ \text{s.t.}\penalty 10000\ \penalty 10000\ R_{i}^{(b)}\cap S(t)\neq\emptyset$
@@ Appendix D Computational complexity
The encoding step requires $\mathcal{O}(n^{2})$ space to store the adjacency matrix. For large datasets, this can be approximated by using a subset of the training data to compute $\mathbf{K}$, at the cost of losing the exact functional equivalence of Eq. 2. Similar tricks are common in the kernel methods literature [100, 107, 5].
@@
[…]
The most efficient decoder is the $k$-NN method, which proceeds in two steps. First, we find the nearest neighbors in embedding space, which for reasonably small $d_{\mathcal{Z}}$ can be completed in $\mathcal{O}(m\log n)$ time using kd-trees. Next, we compute the intersection of all leaf regions for selected neighbors, incurring a cost of $\mathcal{O}(kmd_{\Phi}d_{\mathcal{X}})$.